# m[TODO] #1

### v1

@@ Abstract.
Approximate nearest neighbor search (ANNS) is a fundamental problem in vector databases and AI infrastructures. Recent graph-based ANNS algorithms have achieved high search accuracy with practical efficiency. Despite the advancements, these algorithms still face performance bottlenecks in production, due to the random memory access patterns of graph-based search and the high computational overheads of vector distance. In addition, the performance of a graph-based ANNS algorithm is highly sensitive to parameters, while selecting the optimal parameters is cost-prohibitive, e.g., manual tuning requires repeatedly re-building the index.
This paper introduces VSAG, an open-source framework that aims to enhance the in production performance of graph-based ANNS algorithms. VSAG has been deployed at scale in the services of Ant Group, and it incorporates three key optimizations: (i) efficient memory access: it reduces L3 cache misses with pre-fetching and cache-friendly vector organization; (ii) automated parameter tuning: it automatically selects performance-optimal parameters without requiring index rebuilding; (iii) efficient distance computation: it leverages modern hardware, scalar quantization, and smartly switches to low-precision representation to dramatically reduce the distance computation costs. We evaluate VSAG on real-world datasets. The experimental results show that VSAG achieves the state-of-the-art performance and provides up to 4×4$\times$4 × speedup over HNSWlib (an industry-standard library) while ensuring the same accuracy.
@@
Xiaoyao Zhong, Haotian Li, Jiabao Jin, Mingyu Yang, Deming Chu, Xiangyu Wang, Zhitao Shen, Wei Jia, George Gu, Yi Xie, Xuemin Lin, Heng Tao Shen, Jingkuan Song, Peng Cheng. PVLDB, 14(1): XXX-XXX, 2020.
@@
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/antgroup/vsag.
@@ 1. Introduction
Modern production systems (Wang et al., 2021; Johnson et al., 2019a) typically employ vector quantization to reduce the distance computation cost. Therefore, we set our baseline as HNSW with SQ4 quantization (Zhou et al., 2012). We conduct 1,000 vector queries on the GIST1M [^1] http://corpus-texmex.irisa.fr/. In what follows, we report the performance limitations of graph-based ANNS algorithms, using the experimental evidence of the baseline: (i) high memory access cost: each query needs over 2,959 random vector accesses (total 1.4 MB), causing a 67.42% L3 cache miss rate. The memory-access operations consume 63.02% of the search time. (ii) high parameter tuning cost: if we use the optimal parameters instead of the manually selected values, the QPS can increase from 1,530 to 2,182, by 42.6%. However, brute-force tuning of parameters takes more than 60 hours, which is prohibitively expensive. (iii) high distance computation cost: distance computations still take 26.12% of the search time despite using SQ4 quantization.
@@
    We evaluate algorithms on real datasets in §6. The results show that VSAG can achieve the SOTA performance and outperform HNSWlib by up to 4$\times$ in QPS under the same recall guarantee.
@@ 2. Overview of VSAG Framework
This section provides an overview of our VSAG framework, which includes memory access optimization, automatic parameter tuning, and distance computation optimization. As shown in Figure 1, VSAG integrates these optimizations into different phases of the search process. First, the PRS manages the storage of vector data and graph indices. It stores low-precision quantized codes of vectors alongside high-precision original vectors to support distance computation optimization. This redundant storage is also utilized to balance resource usage and reduce memory access costs. Second, during each hop of the search process when exploring a vector, we employ Deterministic Access Greedy Search to accelerate memory access by minimizing cache misses. Finally, the visited results are pushed into a heap. We apply Selective Re-rank to combine low- and high-precision distances, ensuring effective search performance. The search process then pops the nearest unfinished point from the heap and proceeds to the next hop. Throughout the entire life-cycle of the index, VSAG uses a smart auto-tuner to select parameters that deliver the best search performance across varying environments and retrieval requirements. We will now detail the three main optimizations used in VSAG. The symbols used in this paper are listed in Table 2.
@@ 2.1. Memory Access Optimization
PRS. To better optimize memory access, VSAG introduces a Partial Redundant Storage (PRS) design (see §3.3), which provides a flexible and high-performance storage foundation to optimize both distance computations and memory access while balancing storage and computational resource usage. In production environments constrained by fixed hardware configurations, such as 4C16G (i.e., equipped with 4 CPU cores and 16GB of memory) and 2C8G (i.e., equipped with 2 CPU cores and 8GB of memory), most algorithms frequently exhibit resource utilization imbalances between computational and memory subsystems. During computational processes, CPUs frequently encounter idle cycles caused by cache misses, which hinders their ability to achieve optimal utilization, and thereby limits the system’s QPS.
@@
Balance of Resources. In particular, the system has a parameter called the redundancy ratio $\delta$, providing flexible control over the Balance of Computational Efficiency and Memory Utilization (see §3.3.3). In compute-bound scenarios with high-throughput demands, VSAG adaptively increases the redundancy ratio to mitigate cache contention, thus minimizing CPU idle cycles during memory access while preserving storage efficiency. In contrast, in memory-constrained low-throughput scenarios, the framework strategically reduces redundancy ratio to optimize the index-memory footprint. Under memory-constrained conditions, this optimization enables deployment on reduced instance tiers, thereby curtailing compute wastage.
@@ 3.2.2. Stride Prefetch
Example 0.
3.3. PRS: Flexible Storage Layout Boosting Search Performance
@@
1
2candidate pool $C$leftarrow maximum-heap with size of $ef_{s}$
3
4visited set $V\leftarrow\emptyset$
5
6insert $(x_{i},\tau_{l}(x_{i},x_{q})),\forall x_{i}\in I$ into $C$
7
8while $C$ has un-expended nodes do
@@ 3.3.2. Redundantly Storing Vectors
To address the computational-memory resource imbalance caused by fixed instance specifications (e.g., 4C16G, 2C8G) in industrial applications, we propose PRS to take advantage of hardware prefetching to reduce CPU idle time. A dynamically tunable redundancy ratio $\delta$ controls the proportion of redundantly stored neighbor vectors in graph indices, balancing prefetch efficiency and CPU utilization. When $\delta=1$, full redundancy maximizes hardware prefetch benefits, achieving peak CPU utilization at the cost of higher memory consumption. In contrast, $\delta=0$ eliminates redundancy to minimize memory usage but sacrifices prefetch efficiency. This flexibility enables workload-aware resource optimization as shown in Example 3.5.
@@ 4.1. Parameter Classification
ELPs primarily affect the efficiency (QPS) of the retrieval process and are directly related to the retrieval environment (e.g., prefetch stride $\omega$, prefetch depth $\nu$). Most of these parameters are associated with the execution speed of system instructions rather than the operational flow of the algorithm itself. For instance, prefetch-related parameters mainly influence the timing of asynchronous operations.
@@
ILPs define an index’s core structure and performance. Tuning the construction-related parameters (e.g., $m_{c}$, $\alpha_{c}$) requires rebuilding indices. It is a process far more costly than adjusting QLPs or EIPs. Crucially, ILPs impact both efficiency and effectiveness simultaneously.
@@ 4.2. Search-based Automatic ELPs Tuner
These environment-sensitive parameters exclusively influence algorithmic efficiency metrics (e.g., QPS) while maintaining consistent effectiveness outcomes (e.g., Recall Rate), thereby enabling independent optimization distinct from core algorithmic logic. The VSAG tuning framework implements an optimized three-step procedure:
@@
    Evaluate performance metrics using statistically sampled base vectors.
@@ 4.3. Fine-Grained Automatic QLPs Tuner
To facilitate efficient feature computation, we implement an optimized sorted array structure that replaces conventional heap-based candidate management.
@@ 4.4. Labeling-based Automatic ILPs Tuner
Impact of ILPs. The retrieval efficiency and effectiveness of graph-based indices are fundamentally determined by their structural properties. Modifications to ILPs (e.g., maximum degree $m_{c}$ (Jayaram Subramanya et al., 2019b; Malkov and Yashunin, 2020b), pruning rate $\alpha_{c}$ (Jayaram Subramanya et al., 2019b)) induce concurrent alterations in both graph topology and search dynamics, creating non-linear interactions between retrieval speed and accuracy. The following example demonstrates these interdependent effects:
Example 0.
Reducing $m_{c}$ to 1 causes irreparably damaged graph connectivity. Despite asymptotically increasing the search frontier parameter $ef_{s}$, this configuration achieves near-zero recall accuracy due to path discontinuity. In contrast, when elevating both $m_{c}$ and $\alpha$ to extreme values, the graph degenerates into brute-force search patterns, with each traversal step requiring a full dataset scan, thereby minimizing retrieval throughput.
The necessity for index reconstruction during parameter space exploration creates prohibitive computational costs (typically 2-3 hours per rebuild for million-scale datasets), particularly when optimizing multiple ILPs simultaneously.
@@ 4.4.1. Compression-Based Construction Framework
We empirically observe that indices constructed with varying ILPs exhibit substantial edge overlap. As formalized in Theorem 4.2, when maintaining a fixed pruning rate $\alpha$, a graph built with lower maximum degree $m_{c}$ constitutes a strict subgraph of its higher-degree counterpart.
@@
(Maximum Degree-Induced Subgraph Hierarchy) Let $G^{a}$ and $G^{b}$ denote indices constructed with maximum degrees $m_{c}=a$ and $m_{c}=b$ respectively, under identical initialization conditions and fixed $\alpha$. Given equivalent greedy search outcomes $\mathrm{ANN}_{k}(x)$ for all points $x$ under both configurations, then $a<b\implies G^{a}\subset G^{b}$ where $\forall(x_{i},x_{j})\in E(G^{a}),(x_{i},x_{j})\in E(G^{b})$.
@@
VSAG employs an index compression strategy that constructs the index once using relaxed ILPs while labeling edges with specific ILPs. During the search process, the algorithm dynamically selects edges by leveraging labels, shifting the edge selection process from the index construction phase to the search phase. It enables adaptive edge selection without requiring index reconstruction. Thus, it achieves equivalent search performance to maintaining multiple indices with different parameter configurations, while incurring only the overhead of constructing a single index. The parameter labeling mechanism maintains topological flexibility while enhancing storage efficiency through a compressed index representation. Due to space limitations, please refer to Appendix B for the detailed construction algorithm of VSAG. Then, we illustrate the labeling algorithm, which assigns label to each edge.
@@ 26                  $count\leftarrow count+1$
(Subgraph Inclusion Property with Varying Pruning Rate $\alpha_{c}$) Fix all ILPs except $\alpha_{c}$, and let $G^{a}$ and $G^{b}$ be indices constructed by Algorithm 3 in Appendix using pruning rates $\alpha_{c}=a$ and $\alpha_{c}=b$ respectively, where $a<b$. Suppose that for every data point $x_{i}$, the finite-sized approximate nearest neighbor (ANN) sets $\mathrm{ANN}_{k}(x_{i})$ retrieved during construction remain identical under both $\alpha_{c}$ values. Then $G^{a}$ forms a subgraph of $G^{b}$, i.e., all edges in $G^{a}$ satisfy $(x_{i},x_{j})\in E(G^{b})$.
@@
This theorem reveals a monotonic relationship between pruning rates and graph connectivity. When reducing $\alpha_{c}$, any edge pruned under this stricter parameter setting would necessarily be eliminated under larger $\alpha_{c}$ values. This observation enables us to characterize each edge by its preservation threshold $\alpha_{e}$: the minimal pruning rate required to retain the edge during construction. Consequently, all graph indices constructed with pruning rates $\alpha_{c}\geq\alpha_{e}$ will contain this edge. This threshold-based perspective permits efficient compression of multiple parameterized graph structures into a unified index, where edges are annotated with their respective $\alpha_{e}$ values.
@@
Please refer to Appendix D
for details of tuning ILPs after VSAG is constructed with labels.
@@ 5.1. Distance Computation Cost Analysis
(1) $cost=cost_{lp}+cost_{hp}=n_{lp}\cdot t_{lp}+n_{hp}\cdot t_{hp}$
@@
  • Enhanced quantization precision by parameter optimization (§5.3) mitigates candidate inflation from precision loss, achieving sublinear growth in required low-precision computations ($n_{hp}$).
@@ 5.2. Minimizing Low-Precision Computation Overhead
5.3. Improving Quantization Precision
Scalar quantization (SQ) compresses floating-point vectors by mapping 32-bit float values to lower-bit integer representations in each dimension. In SQ-b (where b denotes bit-width), the dynamic range is uniformly partitioned into $2^{b}$ intervals. For SQ4 (4-bit), this creates 16 intervals over $[0,1]$, where the k-th interval corresponds to $[(k-1)/15,k/15)$. Values within each interval are encoded as their corresponding integer index. However, practical implementations face critical range estimation challenges: direct use of observed min/max values proves suboptimal when data outliers existing.
Example 0.
Figure 6. Truncated scalar quantization.
5.4. Selective Re-rank
6. Experimental Study
6.1. Experimental Setting

### v3

# VSAG: An Optimized Search Framework for Graph-based Approximate Nearest Neighbor Search

@@ Abstract.
Approximate nearest neighbor search (ANNS) is a fundamental problem in vector databases and AI infrastructures. Recent graph-based ANNS algorithms have achieved high search accuracy with practical efficiency. Despite the advancements, these algorithms still face performance bottlenecks in production, due to the random memory access patterns of graph-based search and the high computational overheads of vector distance. In addition, the performance of a graph-based ANNS algorithm is highly sensitive to parameters, while selecting the optimal parameters is cost-prohibitive, e.g., manual tuning requires repeatedly re-building the index. This paper introduces VSAG, an open-source framework that aims to enhance the in production performance of graph-based ANNS algorithms. VSAG has been deployed at scale in the services of Ant Group, and it incorporates three key optimizations: (i) efficient memory access: it reduces L3 cache misses with pre-fetching and cache-friendly vector organization; (ii) automated parameter tuning: it automatically selects performance-optimal parameters without requiring index rebuilding; (iii) efficient distance computation: it leverages modern hardware, scalar quantization, and smartly switches to low-precision representation to dramatically reduce the distance computation costs. We evaluate VSAG on real-world datasets. The experimental results show that VSAG achieves the state-of-the-art performance and provides up to 4×4$\times$4 × speedup over HNSWlib (an industry-standard library) while ensuring the same accuracy.
@@
Xiaoyao Zhong, Haotian Li, Jiabao Jin, Mingyu Yang, Deming Chu, Xiangyu Wang, Zhitao Shen, Wei Jia, George Gu, Yi Xie, Xuemin Lin, Heng Tao Shen, Jingkuan Song, Peng Cheng. PVLDB, 18(12): XXX-XXX, 2025.
@@
Proceedings of the VLDB Endowment, Vol. 18, No. 12 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/antgroup/vsag.
@@ 1. Introduction
Modern production systems (Wang et al., 2021; Johnson et al., 2019a) typically employ vector quantization to reduce the distance computation cost. Therefore, we set our baseline as HNSW with SQ4 quantization (Zhou et al., 2012). We conduct 1,000 vector queries on the GIST1M. In what follows, we report the performance limitations of graph-based ANNS algorithms, using the experimental evidence of the baseline: (i) high memory access cost: each query needs over 2,959 random vector accesses (total 1.4 MB), causing a 67.42% L3 cache miss rate. The memory-access operations consume 63.02% of the search time. (ii) high parameter tuning cost: if we use the optimal parameters instead of the manually selected values, the QPS can increase from 1,530 to 2,182, by 42.6%. However, brute-force tuning of parameters takes more than 60 hours, which is prohibitively expensive. (iii) high distance computation cost: distance computations still take 26.12% of the search time despite using SQ4 quantization.
@@
    We evaluate the algorithms on real datasets with sizes up to 100 million vectors in §6. The results show that VSAG can achieve the SOTA performance and outperform HNSWlib by up to 4$\times$ in QPS under the same recall guarantee.
@@ 2. Overview of VSAG Framework
This section provides an overview of our VSAG framework, which includes memory access optimization, automatic parameter tuning, and distance computation optimization. As shown in Figure 1, VSAG integrates these optimizations into different phases of the search process. First, the PRS manages the storage of vector data and graph indexes. It stores low-precision quantized codes of vectors alongside high-precision original vectors to support distance computation optimization. This redundant storage is also utilized to balance resource usage and reduce memory access costs. Second, during each hop of the search process when exploring a vector, we employ Deterministic Access Greedy Search to accelerate memory access by minimizing cache misses. Finally, the visited results are pushed into a heap. We apply Selective Re-rank to combine low- and high-precision distances, ensuring effective search performance. The search process then pops the nearest unfinished point from the heap and proceeds to the next hop. Throughout the entire life-cycle of the index, VSAG uses a smart auto-tuner to select parameters that deliver the best search performance across varying environments and retrieval requirements. We will now detail the three main optimizations used in VSAG. The symbols used in this paper are listed in Table 2.
@@ 2.1. Memory Access Optimization
PRS. VSAG introduces a Partial Redundant Storage (PRS) design (see §3.3), which provides a flexible and high-performance storage foundation to optimize both distance computations and memory access while balancing storage and computational resource usage. In production environments constrained by fixed hardware configurations, such as 4C16G (i.e., equipped with 4 CPU cores and 16GB of memory) and 2C8G (i.e., equipped with 2 CPU cores and 8GB of memory), most algorithms frequently exhibit resource utilization imbalances between computational and memory subsystems. During computational processes, CPUs frequently encounter idle cycles caused by cache misses, which hinders their ability to achieve optimal utilization, and thereby limits the system’s QPS.
@@
Balance of Resources. In particular, the system uses a parameter called the redundancy ratio $\delta$ to control the Balance of Computational Efficiency and Memory Utilization (see §3.3.3). In compute-bound scenarios with high-throughput demands, VSAG adaptively increases the redundancy ratio to mitigate cache contention, thus minimizing CPU idle cycles during memory access while preserving storage efficiency. In contrast, in memory-constrained low-throughput scenarios, the framework strategically reduces redundancy ratio to optimize the index-memory footprint. Under memory-constrained conditions, this optimization enables deployment on reduced instance tiers, thereby curtailing compute wastage.
@@ 3.2.2. Stride Prefetch
1 candidate pool $C$leftarrow maximum-heap with size of $ef_{s}$
2
3visited set $V\leftarrow\emptyset$
4
5insert $(x_{i},\tau_{l}(x_{i},x_{q})),\forall x_{i}\in I$ into $C$
6
7while $C$ has un-expended nodes do
@@ 13                   $V\leftarrow V\cup\{j\}$
Example 0.
3.3. PRS: Flexible Storage Layout Boosting Search Performance
@@ 3.3.2. Redundantly Storing Vectors
To address the computational-memory resource imbalance caused by fixed instance specifications (e.g., 4C16G, 2C8G) in industrial applications, we propose PRS to take advantage of hardware prefetching to reduce CPU idle time. A dynamically tunable redundancy ratio $\delta$ controls the proportion of redundantly stored neighbor vectors in graph indexes, balancing prefetch efficiency and CPU utilization. When $\delta=1$, full redundancy maximizes hardware prefetch benefits, achieving peak CPU utilization at the cost of higher memory consumption. In contrast, $\delta=0$ eliminates redundancy to minimize memory usage but sacrifices prefetch efficiency. This flexibility enables workload-aware resource optimization as shown in Example 3.5.
@@ 4.1. Parameter Classification
ELPs primarily affect the efficiency (QPS) of the retrieval process and are directly related to the retrieval environment (e.g., prefetch stride $\omega$, prefetch depth $\nu$). Most of these parameters are associated with the execution speed of system instructions rather than the operational flow of the algorithm. For instance, prefetch-related parameters mainly influence the timing of asynchronous operations.
@@
ILPs define an index’s core structure and performance. Tuning the construction-related parameters (e.g., $m_{c}$, $\alpha_{c}$) requires rebuilding indexes. It is a process far more costly than adjusting QLPs or EIPs. Crucially, ILPs impact both efficiency and effectiveness simultaneously.
@@ 4.2. Search-based Automatic ELPs Tuner
These environment-sensitive parameters exclusively influence algorithmic efficiency metrics (e.g., QPS) while maintaining consistent effectiveness outcomes (e.g., Recall), thereby enabling independent optimization distinct from core algorithmic logic. The VSAG tuning framework implements an optimized three-step procedure:
@@
    Evaluate performance metrics using sampled base vectors.
@@ 4.3. Fine-Grained Automatic QLPs Tuner
The candidate size is initially set to a relatively large value (e.g., $ef_{s}$ = 300). During the search process, after traversing certain hops in the graph, we employ the decision tree along with the current features to determine whether the candidate size should be reduced, which effectively prevents overly simple queries from undergoing extensive and unnecessary search process.
@@ 4.4. Labeling-based Automatic ILPs Tuner
Impact of ILPs. The retrieval efficiency and effectiveness of graph-based indexes are fundamentally determined by their structural properties. Modifications to ILPs (e.g., maximum degree $m_{c}$ (Jayaram Subramanya et al., 2019b; Malkov and Yashunin, 2020b), pruning rate $\alpha_{c}$ (Jayaram Subramanya et al., 2019b)) induce concurrent alterations in both graph topology and search dynamics (Yang et al., 2025), creating non-linear interactions between retrieval speed and accuracy. Rebuilding the index while tuning ILPs is computationally expensive(typically taking 2-3 hours per parameter configuration for million-scale datasets).
@@ 4.4.1. Compression-Based Construction Framework
We empirically observe that indexes constructed with varying ILPs exhibit substantial edge overlap. As formalized in Theorem 4.1, when maintaining a fixed pruning rate $\alpha$, a graph built with lower maximum degree $m_{c}$ constitutes a strict subgraph of its higher-degree counterpart.
@@
(Maximum Degree-Induced Subgraph Hierarchy) Let $G^{a}$ and $G^{b}$ denote indexes constructed with maximum degrees $m_{c}=a$ and $m_{c}=b$ respectively, under identical initialization conditions and fixed $\alpha$. Given equivalent greedy search outcomes $\mathrm{ANN}_{k}(x)$ for all points $x$ under both configurations, then $a<b\implies G^{a}\subset G^{b}$ where $\forall(x_{i},x_{j})\in E(G^{a}),(x_{i},x_{j})\in E(G^{b})$.
@@
The lifecycle of the index in VSAG is divided into three distinct phases: building, tuning, and searching. In contrast, traditional indexing frameworks typically consist of only two phases: building and searching. They select edges to construct the index structure using a set of fixed index-level parameters. The reason behind our design is that we believe it is insufficient to rely solely on the raw data for selecting edges during the construction phase. These parameters may need to be dynamically adjusted based on varying user requirements and query complexities.
VSAG employs an index compression strategy that constructs the index once using relaxed ILPs while labeling edges with specific ILPs. Then, VSAG can dynamically select edges by leveraging labels, shifting the edge selection process from the index construction phase to the tuning phase. It enables adaptive edge selection without index reconstruction. Thus, it achieves equivalent search performance to maintaining multiple indexes with different parameter configurations, while incurring only the overhead of constructing a single index. The parameter labeling mechanism maintains topological flexibility while enhancing storage efficiency through a compressed index representation. Due to space limitations, please refer to Appendix B for the detailed construction algorithm of VSAG. Then, we illustrate the labeling algorithm, which assigns label to each edge.
@@ 26                  $count\leftarrow count+1$
(Subgraph Inclusion Property with Varying Pruning Rate $\alpha_{c}$) Fix all ILPs except $\alpha_{c}$, and let $G^{a}$ and $G^{b}$ be indexes constructed by Algorithm 3 in Appendix using pruning rates $\alpha_{c}=a$ and $\alpha_{c}=b$ respectively, where $a<b$. Suppose that for every data point $x_{i}$, the finite-sized approximate nearest neighbor (ANN) sets $\mathrm{ANN}_{k}(x_{i})$ retrieved during construction remain identical under both $\alpha_{c}$ values. Then $G^{a}$ forms a subgraph of $G^{b}$, i.e., all edges in $G^{a}$ satisfy $(x_{i},x_{j})\in E(G^{b})$.
@@
This theorem reveals a monotonic relationship between pruning rates and graph connectivity. When reducing $\alpha_{c}$, any edge pruned under this stricter parameter setting would necessarily be eliminated under larger $\alpha_{c}$ values. Then, we can characterize each edge by its preservation threshold $\alpha_{e}$: the minimal pruning rate required to retain the edge during construction. Consequently, all graph indexes constructed with pruning rates $\alpha_{c}\geq\alpha_{e}$ will contain this edge. This threshold-based perspective permits efficient compression of multiple parameterized graph structures into a unified index, where edges are annotated with their respective $\alpha_{e}$ values.
@@
Please refer to Appendix D for details of tuning ILPs after VSAG is constructed with labels.
@@ 5.1. Distance Computation Cost Analysis
$cost=cost_{lp}+cost_{hp}=n_{lp}\cdot t_{lp}+n_{hp}\cdot t_{hp}$
@@
  • Enhanced quantization precision by parameter optimization (Appendix G) mitigates candidate inflation from precision loss, achieving sublinear growth in required low-precision computations ($n_{hp}$).
@@ 5.2. Minimizing Low-Precision Computation Overhead
ANT-INTERNET 768 9,991,307 1,000 Text
MSMARCO 1024 113,519,750 1,000 Text
5.3. Selective Re-rank
6. Experimental Study
6.1. Experimental Setting